If Ben gives Abe four dollars, Abe will have three times as many dollars as Ben. If Abe gives Ben four dollars, they will have the same amount. How many dollars does each have?
Write a linear system of two equations.
Solve the system by addition or substitution.

Answer by ankor@dixie-net.com(22740)

let a = Abe's amt
let b = Ben's amt
:
If Ben gives Abe four dollars, Abe will have three times as many dollars as Ben.
a + 4 = 3(b-4)
a + 4 = 3b - 12
a = 3b -12 - 4
a = 3b - 16
:
If Abe gives Ben four dollars, they will have the same amount.
a - 4 = b + 4
a = b + 4 + 4
a = b + 8
:
Replace a with (b+8) in the 1st equation
b + 8 = 3b - 16
8 + 16 = 3b - b
24 = 2b
b = $12 has Ben
then obviously
a = $20